I have forgotten how to do 990/N = 45? It has been a long time. Thank you.
Answer by RAY100(1637)

990/N=45,,,,,,multiply both sides by N
.
(990/N)*N = 45 * N
.
990 = 45*N,,,,,,now divide both sides by 45
.
990/45 = (45*N) /45,,,,,,simplify
.
22 = N
.
.
check,,,,990/22 = 45,,,,,ok
.